OpenStudy (anonymous):

solve using substitution?? y=-2x-3 x+y=9

jimthompson5910 (jim_thompson5910):

x+y=9 ... start with the second equation x+(-2x-3)=9 ... replace y with -2x-3 (since y = -2x-3) x-2x-3=9 -x-3 = 9 I'll let you finish up